OpenStudy (anonymous):

Medal and Fan!!!!!!!!!! Which of the following is a polynomial with roots 4, 6, and −7

OpenStudy (anonymous):

y=x^3-3x^2-24x+42 y=x^3-3x^2-46x+168 y=x^3-24x^2-42x+46 y=x^3-24^2-46x+168

OpenStudy (solomonzelman):

A polynomial with roots \(\large\color{blue}{ \displaystyle \rm a }\), \(\large\color{green}{ \displaystyle \rm b }\), and \(\large\color{red}{ \displaystyle \rm c }\) Can be written out as: \(\large\color{black}{ \displaystyle y=(x-\color{blue}{\rm a})(x-\color{green}{\rm b})(x-\color{red}{\rm c}) }\) (then you will expapnd \(\rm \small YOUR \) expression that you get after using this rule.

OpenStudy (solomonzelman):

take it as: 1) your \(\large\color{black}{ \displaystyle \color{blue}{\rm a} }\) is 4 2) your \(\large\color{black}{ \displaystyle \color{green}{\rm b} }\) is 6 3) your \(\large\color{black}{ \displaystyle \color{red}{\rm c} }\) is -7

OpenStudy (anonymous):

the answer is the third one

OpenStudy (solomonzelman):

I doubt that....

OpenStudy (anonymous):

last one

OpenStudy (solomonzelman):

please show your work so that I can tell where the mistake in your expansion is...

OpenStudy (anonymous):

the first one

OpenStudy (lynfran):

(x-4)(x-6)(x+7) (x^2-6x-4x+24)(x+7) (x^2-10x+24)(x+7) x^3+7x^2-10x^2-70x+24x+168...can you simplify it...

OpenStudy (lynfran):

so which one is it?